# Universal Vessel Segmentation for Multi-Modality Retinal Images

### v3

@@ V Conclusions
V-A Limitations
As demonstrated and discussed in the experiment section, our current universal segmentation has several limitations to improve upon. Firstly, although the primary vessels are well predicted, the model tends to over-prune the vessels, which leads to thin vanishing vessels appearing at the peripheral area of the retina or near the macula are missing in the segmentation. This is caused by both the non-neglectable detail loss of the image translation process and the effect of topological loss functions. The effect is prominent in the OCTA and FA modalities where finer vessels are visible. Furthermore, the robustness of the current UVSM to outlier patterns in the peripheral area of UWF images can still be improved.
V-B Conclusion and Future Work
In this paper, we propose the universal retinal vessel segmentation task and the very first algorithm for this task, offering a generalizable and more practical solution to the retinal vessel segmentation problem. Compared with the past literature of retinal vessel segmentation where (1) at most time only Color Fundus modality is studied and; (2) separate models need to be finetuned for different modalities and cameras to achieve good performance, our method is able to perform modality and camera-agnostic vessel segmentation for multi-modal retinal images and achieves comparable performance with SOTA finetuned methods. Although the current model still has limitations and we are working to further improve it, we consider this work as a step forward in the community of retinal image processing. Future works include improving the universal vessel segmentation model to get better performance on fine vessels and to improve the model’s robustness to outlier patterns on the retina.

### v4

@@ V Conclusions
V-A Conclusion
In this paper, we propose the universal retinal vessel segmentation task and the very first method for this task, offering a generalizable and more practical solution to the retinal vessel segmentation problem. Compared with the previous literature on retinal vessel segmentation where (1) in most cases only Color Fundus modality is studied and (2) separate models need to be fine-tuned for different modalities and cameras to achieve good performance, our method allows modality/camera-agnostic vessel segmentation for multi-modal retinal images and achieves comparable performance with SOTA fine-tuned methods.
V-B Limitations and Future Work
As demonstrated and discussed in the experimental section, the current version of our URVSM has several limitations that suggest directions for future improvement.
Firstly, there remains a trade-off between universality and performance. While URVSM performs well on primary vessels, it tends to over-prune finer branches, leading to missed segmentation of thin peripheral vessels or capillaries near the macula. This contributes to a performance gap between URVSM and modality-specific SOTA models. The limitation arises from both the detail loss inherent to the image translation process and the regularizing effect of topological losses.
This issue is especially evident in modalities like FA and OCTA, where capillary-level vessels could be visible. As a result, the current model is not suitable for capillary-level analysis, such as FAZ mapping or OCTA perfusion quantification - which we consider as an important direction for future extensions.
Finally, although the model generalizes well to UWF images overall, its robustness to outlier patterns in the peripheral field of UWF scans can still be improved.